# MOA-2010-BLG-328: Keck and HST Expose the Limits of Occams Razor in Microlensing

### v1

@@ Abstract
We present high resolution follow-up data of the planetary microlensing event MOA-2010-BLG-328, using Keck and the Hubble. Keck data, taken 8 years after the event, reveal a strong lens detection enabling a direct measurement of lens flux and source-lens relative proper motion. We find the relative source-lens proper motion to be $\mu_{\rm rel,Hel}=4.07\pm 0.34\ \rm mas\ yr^{-1}$, with the lens being $\sim 10$ times fainter than the source. The lens was very faint in the Hubble passbands, and the small lens-source separation of $\sim$35 mas made its detection difficult. However, we obtained estimates of the lens magnitudes in Hubble bands by constraining its location to match the Keck K-band detection. The original analysis by Furusawa et al. (2013) reports a degenerate light curve, with several viable models depending on higher-order effects. We attempt to break the degeneracy by remodeling the event using constraints from follow-up data. Our best fit model includes parallax, orbital motion, xallarap and the magnification of a source companion. Models omitting any of these are excluded. However, even with a lens detection the solution remains unclear, as the degeneracy between a nearby late M dwarf and a distant early M dwarf in the disk persists, and cannot be broken with NIR data alone. We conclude the lens is either a $\sim 0.2\ M_{\odot}$ star at $2-3$kpc, or a $\sim 0.5\ M_{\odot}$ star at $4-5$kpc. This study highlights the importance of multi-band data and comprehensive modeling to resolve microlensing degeneracies.
@@ 1 Introduction
There are many different types of techniques that can detect exoplanets (such as radial velocity and stellar transits), however, gravitational microlensing is currently the only one capable of finding planets in wide orbits, beyond the snow line. This parameter space is thought to be the dominant birth place of planets, as predicted by the core accretion theory (Lissauer, 1993; Ida & Lin, 2004; Kennedy et al., 2006). Gravitational microlensing can not only detect ‘cold’ planets down to Earth masses, but can also detect planets that orbit a range of different stars, such as M dwarfs (e.g. OGLE-2005-BLG-071, Udalski et al. (2005); Dong et al. (2009); Bennett et al. (2020)), and even white dwarfs (e.g. MOA-2010-BLG-477, Blackman et al. (2021)). This is because the technique is not reliant on detecting light from the host star. A setback to this technique, however, is that the physical parameters of the system cannot be determined without the addition of higher order effects on the light curve, or direct lens flux and proper motion measurements.
@@
Higher order effects on the light curve can manifest as finite source effects, microlensing parallax, orbital motion of the lens, xallarap, and the magnification of a second source. In the case of MOA-2010-BLG-328 (henceforth, MB10328) Furusawa et al. (2013) they measured parallax, orbital motion and xallarap effects. They considered a “parallax + orbital motion” model, which was degenerate between the close and wide solutions, and they considered a “xallarap” only model. All these models were degenerate, yielding very different results.
@@
In cases such as these, high angular resolution follow-up observations (Bennett et al., 2006, 2007) with either ground based telescopes, like Keck Adaptive Optics, or space telescopes, like the Hubble Space Telescope (HST) can help confirm which solution is correct. Follow-up observations taken several years after the peak magnification can directly measure the lens-source relative proper motion in addition to the lens flux. Once these are added as constraints to the light curve model, several mass-distance relations can be obtained, and thus the physical properties of the lensing system can be determined (e.g. Beaulieu (2018); Bhattacharya et al. (2018); Bennett et al. (2020); Vandorou et al. (2020); Terry et al. (2021, 2022)).
@@
In this paper we present high resolution adaptive optics (AO) observations from Keck-II’s NIRC2 camera and the Hubble Space Telescope (HST) where we obtain constraints on the source-lens relative proper motion, and individual magnitudes, concluding that this is a binary lens binary source microlensing system. Using these constraints, we reanalyze the light curve and find that incorporating additional higher-order effects is necessary beyond what was previously considered in Furusawa et al. (2013). We therefore model the light curve with finite-source effects, microlenisng parallax, orbital motion, xallarap, and magnification of the second source. This challenges the usual assumption that the “simplest explanation is the best”, a philosophical principle known as Occam’s Razor. Our “challenge” of it implies that this event presents a complex and degenerate set of solutions where multiple higher order effects must be considered. Keck and HST provide crucial follow-up data, but do not fully resolve the degeneracy, highlighting fundamental challenges in microlensing analysis.
@@
The paper is organised as follows; in Section 2 we present the microlensing event MB10328 with results from the detection paper by Furusawa et al. (2013) and the challenges they faced. In Section 3 we present new high resolution follow-up data that was acquired in 2018 from Keck and HST, and the analysis of this data. The results of this data are subsequently used as constraints for the light curve remodeling, and presented in Section 4, where we also explore the various higher order effects and degeneracies. The physical parameters of the lens system are described in Section 5. Finally, we discuss the overall results and conclude the paper in Section 6.
@@ 2 Microlensing event MOA-2010-BLG-328
The microlensing event MB10328 was first detected by the Microlensing Observations in Astrophysics (MOA) group on 16 June 2010. The event is located at (R.A., Dec) = (17:57:59.12, –30:42:54.63) and Galactic coordinates $(l,b)=(-0.16,-3.21)$ degrees. An anomaly in the predicted single lens light curve model was observed, and an alert was issued to other microlensing groups on 27 July 2010. Shortly after other follow-up groups began to observe the event, and data was obtained from the Microlensing Follow-Up Network ($\mu$Fun; Gould et al. (2006)), the Probing Lensing Anomalies NETwork (PLANET; Beaulieu et al. (2006)), Microlensing Network for the Detection of Small Terrestrial Exoplanets (MiNDSTEp; Dominik et al. (2010)) and RoboNet (Tsapras et al., 2009).
The data, modelling and analyses are presented in Furusawa et al. (2013). Their light curve model revealed the effects of a microlensing parallax due to the orbital motion of the Earth (Smith et al., 2003) and the orbital motion of the source star, which is often referred to as the xallarap effect. Although they do not completely exclude the xallarap solution, the data prefers the parallax plus orbital motion solution by $\Delta\chi^{2}=5$. At such a low level of significance this could be influenced by systematics, which often affects parallax solutions in microlensing events. For events that have a relatively weak microlensing parallax, approximate symmetry exists in the plane where the lens is replaced by its mirror image. That symmetry is broken for this event however, and the $u_{0}<0$ and $u_{0}>0$ solutions differ by $\Delta\chi^{2}=78$, where $u_{0}$ is the impact parameter in units of Einstein radii. The light curve parameters for each model are presented in Table 4.
The source color and magnitude were derived from CTIO V and I band photometry, and the observed I band magnitude of the source star from the model and OGLE data. In combination with the color-color relations from Bessell & Brett (1988) they found the $V$, $I$ and $K$ magnitudes of the source for the parallax plus orbital motion solutions and the xallarap solution. These are presented in Table 1.
@@
From these source colors and magnitudes they calculated the angular Einstein ring radius,$\theta_{E}$, and relative source-lens proper motion, $\mu$. Combining these with a measured parallax, they obtained physical parameters for the system for the close and wide parallax plus orbital motion model, assuming a source distance $D_{\rm S}=8.0\pm 0.3$. They found a system with lens mass of $M_{\rm L}\sim 0.1\ M_{\odot}$ with a super Earth planetary companion. For the xallarap model, the probability distributions of the physical parameters were estimated using a Bayesian analysis. These results indicated a $M_{\rm L}\sim 0.6\ M_{\odot}$ lens mass with a Saturn-mass planetary companion. See Table 5 for more details.
Using high resolution follow-up observations with Keck and HST we directly measure the source and lens flux, in addition to the relative proper motion (direction and magnitude). With this data as constraints we attempt to break the degenaracies in the light curve model and find a plausible solution for this system.
@@ 3.1 Keck NIRC2
On August 5 and 6 2018, MB10328 was observed again, using Keck II’s NIRC2 narrow camera which has a plate scale of 9.94 mas/pixel. The images had an exposure time of 30 seconds in the $K$ band. 10 images were obtained on August 5, and 30 on August 6. The images were reduced using ‘KAI’, which is a data-reduction pipeline specifically designed to reduce Keck AO data (Lu, 2022). The pipeline includes the standard dark current correction, flat-fielding and sky correction, in addition to bad pixel and cosmic ray masking. The images are combined using a Jackknife routine (Quenouille (1949, 1956); Tukey (1958)) where $\rm N-1$ frames are stacked, where N is the number of total frames. Therefore, we produce 10 different stacks of 9 frames for the August 5 observations, and 30 different stacks of 29 frames for the August 6 data, since one frame is removed each time. The target star can be seen in Figure 1 for both nights. This process was repeated for both August 5 and 6, and the subsequent analyses on the images was conducted separately, and for the final result we use the average of these two sets of data. Using this Jackknife stacking method is standard procedure now when analysing high resolution follow-up images with Keck, with further details to be found in Terry et al. (2021); Bhattacharya et al. (2021).
@@
3.1.1 PSF Fitting Photometry &\&& Astrometry
@@
We use a modified version of DAOPHOT-II (Stetson, 1987) called DAOPHOT-MCMC (Terry et al., 2021) which runs Markov Chain Monte Carlo (MCMC) sampling on the target star. The first step to this routine is to build an empirical PSF model using reference stars of similar brightness and PSF shape in the target’s field. This is done for each individual jackknife frame, as the PSF shape can vary from image to image. We select reference stars with magnitude ranges of $-0.3<m<0.3$ and separations $-3^{\prime\prime}<r<3^{\prime\prime}$ from the target.
@@
The main source of uncertainty in our data stems from the PSF variations between the individual Keck frames, and across each frame. Using the jackknife method to stack our frames allows us to determine this uncertainty between the images. The errors are estimated by the $\sqrt{N-1}$ multiplied by the root-mean-square of the best fit parameters from each $N-1$ combinations as shown in the equation below:
@@ 3.1.3 Source-lens relative proper motion and flux ratio
$\mu_{\rm Hel,East}$ ($\rm mas\ yr^{-1}$) $0.708\pm 0.103$
$\mu_{\rm Hel,North}$ ($\rm mas\ yr^{-1}$) $4.371\pm 0.104$
Lens / Source flux $0.120\pm 0.003$
@@
By directly measuring the separating between the source and lens positions in the Keck data, and knowing that 7.93 years have passed since peak magnification, we calculate the heliocentric relative proper motion to be $(\mu_{\rm Hel,E},\mu_{\rm Hel,N})=(0.708,4.371)\pm(0.103,0.104)\ \rm mas\ yr^{-1}$, where E and N denote the East and North direction, respectively. Therefore the amplitude of the relative proper motion vector is $\mu_{\rm rel,1Hel}=4.428\pm 0.104\ \rm mas\ yr^{-1}$.
@@
We can compare this source $K$ band value with those found by Furusawa et al. (2013) (see Table 1), and we find that our measured Keck value is $\sim 0.8$ mag brighter. This is unsurprising since the system likely has a binary source. Therefore, when remodeling the light curve in Section 4, a binary lens binary source scenario is explored which includes the higher order effects of parallax, orbital motion and xallarap.
@@ 3.2 HST WFC3/UVIS
The microlensing event MB10328 was observed twice with the Hubble Space Telescope’s (HST) WFC3/UVIS camera. The first set of observations were taken 27 May 2018 in the F814W and F [^5] W filters as part of the GO-15455 proposal (PI: Bennett).
All data was obtained from the Mikulski Archive for Space Telescopes (MAST) and flat-fielded, stacked and corrected for disortions. Pixel area distortion, astrometry and PSF photometry were carried out on the F814W and F [^5] W data separately with a modified version of the code used by Bennett et al. (2015) and Bhattacharya et al. (2018). This code does not use any resampling, and instead analyses the data from the original images in order to mitigate any loss in resolution that a combined, dithered, undersampled image would. In addition, due to HST’s relative smaller aperture and large undersampled pixels, its images have worse angular resolution compared to Keck. However, HST compensates for this by having a much more stable PSF.
@@
We fit a single star PSF model to produce the 1-star residual (see Figure 4), and then a two star PSF model. The single star residual model does not visually show the signal that we expect from two blended stars (as seen in the Keck images). However, from the Keck images we have a clear result on the position of the lens relative to the source. Even though we cannot “see by eye” a residual signal, we run a constrained fit using an MCMC routine on the position and flux of the lens and source, using the constraints on $\mu_{\rm Hel}$ from the Keck 2018 data. We calibrate the HST data to the OGLE-III catalog (Szymański et al., 2011) by matching bright isolated stars. Therefore we measure the V and I magnitudes of the source (primary and secondary source total) and the lens which has a known position relative to the source. These values are shown in Table 3. For the combined total flux of the blended target (lens and source) in HST we find:
Figure 4: From left to right: first panel shows the stacked image of the 2018 HST $I$-band data, the second shows a zoom of the target with the constrained lens and source positions indicated in cyan and red, respectively. The third and fourth panels show the 1 and 2 star fits to the target, respectively. The lens is not visible in these images.
@@
The 1-star model yields a $\chi^{2}=510.6262$ compared to that of the 2-star model with $\chi^{2}=510.8950$. Although the 1-star model yields a higher $\chi^{2}$, the difference is very small, suggesting that the results of the 2-star model can also be reasonable. Nevertheless, going forward we only use the total color and magnitude of the source and lens as a constraint in the image constrained light curve model (see Section 4). The lens is faint in the $V$ and $I$ passbands, and the lens-source separation is small compared to the HST FWHM. Therefore, getting a good measurement of both the lens-source flux ratio and their separation is difficult. However, if the lens-source separation is known, then the lens flux can be measured in a constrained fit Bennett et al. (2007). This constrained fit yields a measurement of the lens $I$ band magnitude and a very marginal measurement of the $V$ band magnitude.
@@
4 Image Constrained Light Curve Models
Table 4: Light Curve Model Parameters using the standard solution, the microlensing parallax plus orbital motion close and wide solution, and the xallarap solution. Parameter discriptions can be found throughout Section 2.
      Parallax plus orbital motion
      Parameter       ($u_{0}<0$)       ($u_{0}>0$)       Xallarap       New model
      $t_{\rm E}$ (days)       $62.6\pm 0.6$       $64.2\pm 0.6$       $61.8\pm 0.3$       $62.64\pm 0.52$
      $t_{0}$ (${\rm HJD}^{\prime})$       $5378.683\pm 0.014$       $5378.694\pm 0.017$       $5378.706\pm 0.013$       5378.211$\pm$ 0.506
      $t_{*}$ (days)       0.0582 $\pm$ 0.0063       $0$$\pm$ 0.0109       0.0834 $\pm$ 0.0007       0.0804 $\pm$ 0.0177
      $u_{0}$       $-0.0721\pm 0.0008$       $0.0716\pm 0.0007$       $-0.0741\pm 0.0004$       -0.0571$\pm$ 0.0091
      s       $1.154\pm 0.016$       $1.180\pm 0.028$       $1.220\pm 0.012$       1.224 $\pm$ 0.005
      $\pi_{\rm E,N}$       1.01 $\pm$ 0.06       0.72 $\pm$ 0.05       –       0.290 $\pm$ 0.140
      $\pi_{\rm E,E}$       -0.51 $\pm$ 0.04       -0.39 $\pm$ 0.03       –       -0.154 $\pm$ 0.134
      $ds/dt\ (\rm day^{-1})$       $0.00251\pm 0.00063$       $0.00141\pm 0.00116$       –       $0.000016\pm 0.000027$
      $\omega\ \rm(rad/day)$       $-0.00739\pm 0.00039$       $-0.00139\pm 0.00060$       –       $-0.000253\pm 0.000415$
      $t_{0,2}$ (HJD’)       –       –       –       5333.214$\pm$ 12.271
      $u_{0,2}$       –       –       –       0.548$\pm$ 0.959
      f2rI       –       –       –       0.309 $\pm$ 0.0433
      f2rV       –       –       –       0.443$\pm$ 0.098
      $\rm dt_{\rm E21}$ (days)       –       –       –       5.418$\pm$ 7.547
      $\rm d\theta$ (mas)       –       –       –       0.36799 $\pm$ 0.05933
      $t_{\rm fix}$ (days)       -       -       -       5406.0
      $D_{\rm S}$ (kpc)       –       –       7.637 $\pm$ 0.581
      $\chi^{2}$       5657.75       5660.31       5652.59       4654.9
      $dof$       5660       5660       5658       4682
To determine the lens system’s physical parameters such as mass and distance, we still need properties from the microlensing light curve model. In cases where the light curve model is degenerate, we can use constraints placed on $\mu_{\rm rel}$ and the lens flux and color to try and break the degeneracy. A collection of model parameters can be generated using MCMC calculations that are consistent with the light curve and models that are then inconsistent with constraints from the high resoltion follow-up data are excluded.
To determine the mass of the lens system from the relative proper-motion constraint, we need to know the distance to the source star ($D_{\rm S}$) as shown in equation 4. We include $D_{\rm S}$ as a light curve model parameter and use the Koshimoto et al. (2021) Galactic model as a prior. This prior is used to weight the entries in a sum of Markov chain values, not directly in the light curve modeling. Therefore, we find that the source is at $D_{\rm S}=7.771\pm 0.669$ kpc.
The modeling for MB10328 is particularly challenging as it includes many higher order effects, which are not typically modeled simultaneously as seen in Furusawa et al. (2013). However, in the remodeling of this event we include all higher order effects; microlensing parallax, orbital motion, xallarap and magnification of the second source. The modeling is done with a modified version of the light curve modeling code eesunhong [^1] https://github.com/golmschenk/eesunhong (Bennett & Rhie, 1996; Bennett et al., 2010). The eesunhong code follows an image-centered ray shooting method and includes constraints on both the brightness and relative proper motion of the source and lens from the high resolution follow-up data. For a full discreption of this methodology we refer the reader to Bennett et al. (2024). To speed up the light curve modeling, most of the modeling was done with photometry data that was binned to an interval of 0.01 days, after verifying that this binning had no significant effect on the model parameters.

### v2

@@ Abstract
We present high-resolution follow-up data of the microlensing event MOA-2010-BLG-328, using Keck and the Hubble Space Telescope. Keck data, taken 8 years after the event, reveal a strong lens detection enabling measurement of lens flux and source-lens relative proper motion. We find $\mu_{\rm rel,Hel}=4.07\pm 0.34\ \rm mas\ yr^{-1}$, with the lens $\sim 10$ times fainter than the source. The lens was very faint in the Hubble passbands, and the small $\sim$35 mas lens-source separation made its detection difficult. However, we estimated lens magnitudes in Hubble bands by constraining its location to match the Keck $K$-band detection. Furusawa et al. (2013) reported a degenerate light curve, with viable models including microlensing parallax and lens orbital motion, or xallarap. We attempt to break this degeneracy by remodeling the event with constraints from high-resolution imaging, but find none of their models fit. Instead, fitting the follow-up data requires microlensing parallax, lens orbital motion, and a magnified binary companion to the host, with xallarap. The high-resolution data do break the “ecliptic degeneracy” common in Galactic bulge microlensing events with parallax signals. Models omitting any of these effects are excluded. However, a new degeneracy allows different combinations of parallax, orbital motion, xallarap, and source companion magnification to explain the light curve. These require either a nearby late M dwarf host or a more distant early M dwarf host, indistinguishable with NIR data alone.
@@ 1 Introduction
There are many different types of techniques that can detect exoplanets (such as radial velocity and stellar transits), however, gravitational microlensing is currently the only one capable of finding low-mass planets in wide orbits, beyond the snow line. The direct imaging technique can also detect planets in wide-orbits, however, they are limited to more massive planets ($1\ M_{\rm jup}$). This wide-orbit parameter space is thought to be the dominant birth place of planets, as predicted by the core accretion theory (Lissauer, 1993; Ida & Lin, 2004; Kennedy et al., 2006). Gravitational microlensing cannot only detect ‘cold’ planets down to Earth masses, but can also detect planets that orbit a range of different stars, such as M dwarfs (e.g. OGLE-2005-BLG-071, Udalski et al. (2005); Dong et al. (2009); Bennett et al. (2020)), and even white dwarfs (e.g. MOA-2010-BLG-477, Blackman et al. (2021)). This is because the technique is not reliant on detecting light from the host star. A setback to this technique, however, is that the physical parameters of the system cannot be determined without the addition of higher order effects on the light-curve, or direct lens flux and proper motion measurements.
@@
Higher order effects on the light-curve can manifest as finite source effects, microlensing parallax, orbital motion of the lens, xallarap, and the magnification of a second source. In the case of MOA-2010-BLG-328 (henceforth, MB10328) Furusawa et al. (2013) measured parallax, orbital motion and xallarap effects. They considered a “parallax + orbital motion” model, which was degenerate between the close and wide solutions, and they considered a “xallarap” only model. All these models were degenerate, yielding very different results.
@@
In cases such as these, high angular resolution follow-up observations (Bennett et al., 2006, 2007) with either ground based telescopes, like Keck Adaptive Optics, or space telescopes, like the Hubble Space Telescope (HST) can help confirm which solution is correct. Follow-up observations taken several years after the peak magnification can directly measure the lens-source relative proper motion in addition to the lens flux. Once these are added as constraints to the light-curve model, several mass-distance relations can be obtained, and thus the physical properties of the lensing system can be determined (e.g. Beaulieu (2018); Bhattacharya et al. (2018); Bennett et al. (2020); Vandorou et al. (2020); Terry et al. (2021, 2022)).
@@
In this paper we present high resolution adaptive optics (AO) observations from Keck-II’s NIRC2 camera and the Hubble Space Telescope (HST) where we obtain constraints on the source-lens relative proper motion, and individual magnitudes, concluding that this is a binary lens binary source microlensing system. Using these constraints, we reanalyze the light-curve and find that incorporating additional higher-order effects is necessary beyond what was previously considered in Furusawa et al. (2013). We therefore model the light-curve with finite-source effects, microlenisng parallax, orbital motion, xallarap, and magnification of the second source. This challenges the usual assumption that the “simplest explanation is the best”, a philosophical principle known as Occam’s Razor. Our “challenge” of it implies that this event presents a complex and degenerate set of solutions where multiple higher order effects must be considered. Keck and HST provide crucial follow-up data, but do not fully resolve the degeneracy, highlighting fundamental challenges in microlensing analysis.
@@
The paper is organised as follows; in Section 2 we present the microlensing event MB10328 with results from the detection paper by Furusawa et al. (2013) and the challenges they faced. In Section 3 we present new high resolution follow-up data that was acquired in 2018 from Keck and HST, and the analysis of this data. The results of this data are subsequently used as constraints for the light-curve remodeling, and presented in Section 4, where we also explore the various higher order effects and degeneracies. The physical parameters of the lens system are described in Section 5. Finally, we discuss the overall results and conclude the paper in Section 6.
@@ 2 Microlensing event MOA-2010-BLG-328
[…]
The data, modelling and analyses are presented in Furusawa et al. (2013). Their light-curve model revealed the effects of a microlensing parallax due to the orbital motion of the Earth (Smith et al., 2003) and the orbital motion of the source star, which is often referred to as the xallarap effect. Although they do not completely exclude the xallarap solution, the data prefers the parallax plus orbital motion solution by $\Delta\chi^{2}=5$. At such a low level of significance this could be influenced by systematics, which often affects parallax solutions in microlensing events. For events that have a relatively weak microlensing parallax, approximate symmetry exists in the plane where the lens is replaced by its mirror image. That symmetry is broken for this event however, and the $u_{0}<0$ and $u_{0}>0$ solutions differ by $\Delta\chi^{2}=78$, where $u_{0}$ is the impact parameter in units of Einstein radii.
The source color and magnitude were derived from CTIO V and $I$ band photometry, and the observed I band magnitude of the source star from the model and OGLE data. In combination with the color-color relations from Bessell & Brett (1988) they found the $V$, II and $K$ magnitudes of the source for the parallax plus orbital motion solutions and the xallarap solution. These are presented in Table 1.
@@
From these source colors and magnitudes they calculated the angular Einstein ring radius, $\theta_{E}$, and relative source-lens proper motion, $\mu$. Combining these with a measured parallax, they obtained physical parameters for the system for the close and wide parallax plus orbital motion model, assuming a source distance $D_{\rm S}=8.0\pm 0.3$. They found a system with lens mass of $M_{\rm L}\sim 0.1\ M_{\odot}$ with a super Earth planetary companion. For the xallarap model, the probability distributions of the physical parameters were estimated using a Bayesian analysis. These results indicated a $M_{\rm L}\sim 0.6\ M_{\odot}$ lens mass with a Saturn-mass planetary companion.
MB10328 was reobserved with Keck and HST, and the target was identified in both using the initial MOA survey data as seen in Figure 1. A sample of 624 difference images were used where the event was undergoind lensing, and the individual centroid positions on this sample were measured with the optomised PSF used for the light-curve photometry. The scatter of this measurement determined the error box boundaries in the MOA frame, with the width being $\sim 0.1$ arcseconds. The MOA frame was matched to the HST frame using an affine transorm relation, derived by cross referencing the same stars in both frames. This resulted in a total of 44 matching star-pair positions, which were used to derive the transformation coefficients. The MOA-HST transformation was then used to transform the MOA error box and the clustering of individual measurements around the star. The MOA and Keck frames were matched in a similar process, finding 50 matching star-pairs. Transforming the MOA measurements onto the Keck frame resulted in the same star being identified as in the HST images.
Using high resolution follow-up observations with Keck and HST we directly measure the source and lens flux, in addition to the relative proper motion (direction and magnitude). With this data as constraints we attempt to break the degenaracies in the light-curve model and find a plausible solution for this system.
[…]
@@ 3.1 Keck NIRC2
On August 5 and 6 2018, MB10328 was observed again, using Keck II’s NIRC2 narrow camera which has a plate scale of 9.94 mas/pixel. The images had an exposure time of 30 seconds in the $K$ band. 10 images were obtained on August 5, and 30 on August 6. The images were reduced using ‘KAI’, which is a data-reduction pipeline specifically designed to reduce Keck AO data (Lyke et al., 2017; Lockhart et al., 2019). The pipeline includes the standard dark current correction, flat-fielding and sky correction, in addition to bad pixel and cosmic ray masking. The images are combined using a Jackknife routine (Quenouille (1949, 1956); Tukey (1958)) where $\rm N-1$ frames are stacked, where N is the number of total frames. Therefore, we produce 10 different stacks of 9 frames for the August 5 observations, and 30 different stacks of 29 frames for the August 6 data, since one frame is removed each time. The target star can be seen in Figure 2 for both nights. This process was repeated for both August 5 and 6, and the subsequent analyses on the images was conducted separately, and for the final result we use the average of these two sets of data (see Table 2). Using this Jackknife stacking method is standard procedure now when analysing high resolution follow-up images with Keck, with further details to be found in Terry et al. (2021); Bhattacharya et al. (2021).
@@
3.1.1 PSF Fitting Photometry &\& Astrometry
@@
We use a modified version of DAOPHOT-II (Stetson, 1987) called DAOPHOT-MCMC (Terry et al., 2021) which runs Markov Chain Monte Carlo (MCMC) sampling on the target star. The first step to this routine is to build an empirical PSF model using reference stars of similar brightness and PSF shape in the target’s field. This is done for each individual jackknife frame, as the PSF shape can vary from image to image. We select reference stars with magnitude ranges of 0.3<m<0.3-0.3<m<0.3 and separations $3^{\prime\prime}<r<3^{\prime\prime}$ from the target.
@@
The main source of uncertainty in our data stems from the PSF variations between the individual Keck frames, and across each frame. Using the jackknife method to stack our frames allows us to determine this uncertainty between the images. The errors are estimated by the $\sqrt{N-1}$ multiplied by the root-mean-square of the best fit parameters from each N1N-1 combinations as shown in the equation below:
@@ 3.1.3 Source-lens relative proper motion and flux ratio
$\mu_{\rm rel,Hel,East}$ ($\rm mas\ yr^{-1}$) $0.708\pm 0.103$
$\mu_{\rm rel,Hel,North}$ ($\rm mas\ yr^{-1}$) $4.371\pm 0.104$
Lens / Source flux $0.119\pm 0.003$
@@
By directly measuring the separation between the source and lens positions in the Keck data, and knowing that 7.93 years have passed since peak magnification, we calculate the heliocentric relative proper motion to be $(\mu_{\rm rel,Hel,E},\mu_{\rm rel,Hel,N})=(0.708,4.371)\pm(0.103,0.104)\ \rm mas\ yr^{-1}$, where E and N denote the East and North direction, respectively. Therefore the amplitude of the relative proper motion vector is $\mu_{\rm rel,1Hel}=4.428\pm 0.104\ \rm mas\ yr^{-1}$.
@@
We can compare this source $K$ band value with those found by Furusawa et al. (2013) (see Table 1), and we find that our measured Keck value is $\sim 0.8$ mag brighter. This is unsurprising since the system likely has a binary source. Therefore, when remodeling the light-curve in Section 4, a binary lens binary source scenario is explored which includes the higher order effects of parallax, orbital motion and xallarap.
@@ 3.2 HST WFC3/UVIS
The microlensing event MB10328 was observed twice with the Hubble Space Telescope’s (HST) WFC3/UVIS camera. The first set of observations were taken 27 May 2018 in the F814W and F [^5] W filters as part of the GO-15455 proposal (PI: Bennett).  [^1] The HST data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/yftx-fp58 (catalog doi: 10.17909/yftx-fp58).
All data were flat-fielded, stacked and corrected for disortions. Pixel area distortion, astrometry and PSF photometry were carried out on the F814W and F [^5] W data separately with a modified version of the code used by Bennett et al. (2015) and Bhattacharya et al. (2018). This code does not use any resampling, and instead analyses the data from the original images in order to mitigate any loss in resolution that a combined, dithered, undersampled image would. In addition, due to HST’s relative smaller aperture and large undersampled pixels, its images have worse angular resolution compared to Keck. However, HST compensates for this by having a much more stable PSF.
@@
We fit a single star PSF model to produce the 1-star residual (see Figure 5), and then a two star PSF model. The single star residual model does not visually show the signal that we expect from two blended stars (as seen in the Keck images). However, from the Keck images we have a clear result on the position of the lens relative to the source. Even though we cannot “see by eye” a residual signal, we run a constrained fit using an MCMC routine on the position and flux of the lens and source, using the constraints on $\mu_{\rm rel}$ from the Keck 2018 data. We calibrate the HST data to the OGLE-III catalog (Szymański et al., 2011) by matching bright isolated stars. Therefore we measure the V and I magnitudes of the source (primary and secondary source total) and the lens which has a known position relative to the source. These values are shown in Table 3. For the combined total flux of the blended target (lens and source) in HST we find:
Figure 5: From left to right: first panel shows the stacked image of the 2018 HST II-band data, the second shows a zoom of the target with the constrained lens and source positions indicated in cyan and red, respectively. The third and fourth panels show the 1 and 2 star fits to the target, respectively. The faint lens star cannot be directly identified in these images, however, since we know the lens-source separation from the Keck analysis, we can measure the lens and source magnitudes with a fit that constrains their separation to be the same as measured from the Keck data.
@@
The 1-star model yields a $\chi^{2}=510.6262$ compared to that of the 2-star model with $\chi^{2}=510.8950$. Although the 1-star model yields a higher $\chi^{2}$, the difference is very small, suggesting that the results of the 2-star model can also be reasonable. Nevertheless, going forward we only use the total color and magnitude of the source and lens as a constraint in the image constrained light-curve model (see Section 4). The lens is faint in the $V$ and $I$ passbands, and the lens-source separation is small compared to the HST FWHM. Therefore, getting a good measurement of both the lens-source flux ratio and their separation is difficult. However, if the lens-source separation is known, then the lens flux can be measured in a constrained fit Bennett et al. (2007). This constrained fit yields a measurement of the lens $I$ band magnitude and a very marginal measurement of the $V$ band magnitude.
@@
4 Image Constrained Light-Curve Model
To obtain a full solution of the lens system’s physical properties (lens mass, distance, planet mass, and projected separation), a combination of light-curve model and high resolution follow-up parameters is needed. This process requires that we first remodel the light-curve using constraints on $\mu_{\rm rel}$, the lens and source flux, and color from the high resolution data as described in Bennett et al. (2024). The MCMC results from this modeling procedure are therefore consistent with the high resolution follow-up data. Models that are consistent with the light-curve data, but not the high angular resolution follow-up data are excluded by the large $\chi^{2}$ contributions from the high resolution data constraints. Then, the lens system parameters can be determined by summing over the light-curve model MCMC results using a Galactic prior, in combination with mass-distance and mass-luminosity relations. We discuss these relations in Section 5.
With the measured relative proper motion magntitude and direction, $\mu_{\rm rel}$ from Keck and its uncertainty, we can constrain the Einstein ring radius $\theta_{\rm E}$, and the direction of the microlensing parallax $\pi_{\rm E}$, because the vector $\pi_{\rm E}$ is parallel to the $\mu_{\rm rel}$ vector once it has been converted from heliocentric to geocentric reference frame (see Section 5 and equations therein). We also constrain the lens star brightness using the measured lens $K$ flux from Keck, and the combined source and lens $I$ and $V$ flux from our HST measurements.
Figure 6: Best fit 2L2S light-curve for MB10328 using constraints from the Keck and HST high resolution follow-up data. The middle panel is a close-up of the upper pannel, indicating the planetary anomally. The lower plot shows the residuals. The parameters are described in Section 4 and Table 4.
[…]